the present ratio between the age of father and son is 20:1.
After 6 years the ratio will become 5:1.
find the present age of father
A,18 YEARS B, 24 YEARS C,32 YEARS D,35 YEARS

Answer by jim_thompson5910(35256)

x = age of father
y = age of son

Present ratio of father and son = 20:1
so
(age of father)/(age of son) = 20/1
x/y = 20/1
x/y = 20
x = 20*y ... multiply both sides by y

"after 6 years, the ratio will become 5:1", so,
(x+6)/(y+6) = 5/1
(x+6)/(y+6) = 5
x+6 = 5*(y+6)
20y+6 = 5*(y+6) ... replace x with 20y (see above)
20y+6 = 5y+30
20y+6 - 5y = 5y+30-5y ... Subtract 5y from both sides.
15y+6 = 30
15y+6-6 = 30-6 ... Subtract 6 from both sides.
15y = 24
15y/15 = 24/15 ... Divide both sides by 15.
y = 24/15
y = 8/5
y = 1.6

Use this value of y to find x
x = 20*y
x = 20*1.6 ... replace y with 1.6
x = 32

The father is currently 32 years old